A rectangle and a square have the same area. The length of the rectangle is 168 feet more than 2 times its width. The Length of a side of a square is 72 feet. What are the dimensions of the rectangle?
Answer by lwsshak3(11628)

A rectangle and a square have the same area. The length of the rectangle is 168 feet more than 2 times its width. The Length of a side of a square is 72 feet. What are the dimensions of the rectangle?
***
let x=width of rectangle
2x+168=length of rectangle
Area of rectangle=area of square
x(2x+168)=72^2
2x^2+168x=5184
2x^2+168x-5184=0
x^2+84x-2592=0
(x-24)(x+108)=0
x=-108 (reject)
or
x=24
2x+168=48+168=216
dimensions of the rectangle: 24 ft by 216 ft